valkraybff:

The scale model of a rectangular garden is 1.5 ft by 4 ft. The scale model is enlarged by a scale factor of 7 to create the actual garden. What is the area of the actual garden? 6 Feet squared 42 Feet squared 252 Feet squared 294 Feet squared

Extrinix:

So just like the other one we did, you multiply your scale factor by your original measurements. \(\color{lime}{1.5 \times 4}\) -> Your original area Then you can multiply the scale factor. \(\color{lime}{(1.5 \times 7) \times (4 \times 7)}\) -> Your new area

Extrinix:

The way it works is pretty simple, if your scale factor is a fraction, you're "decreasing" the area, and if it's a whole number, you're "increasing" your area.

valkraybff:

can you dumb it down

Extrinix:

\(\color{lime}{(1.5 \times 7) \times (4 \times 7)}\) Plain and simple

valkraybff:

=

Extrinix:

Can't give a direct answer, you have to solve that yourself.

CrazyHotGhost:

@valkraybff wrote:
can you dumb it down
NAHHHHHHHHHHH 😭😭

caydenmorrison:

@valkraybff wrote:
can you dumb it down
How x.x

CrazyHotGhost:

@caydenmorrison wrote:
@valkraybff wrote:
can you dumb it down
How x.x
righttttttttt

valkraybff:

dude you guy need some manner

CrazyHotGhost:

no you need some brain cells

CrazyHotGhost:

like damn my brain is fried and i understand this better than you do.

xXDerpyPugXx:

@extrinix wrote:
\(\color{lime}{(1.5 \times 7) \times (4 \times 7)}\) Plain and simple
Just use a calculator

valkraybff:

dude talk about some other girl because I am smart just need help like evryone hear

Extrinix:

There's a calculator on here

CrazyHotGhost:

@valkraybff wrote:
dude talk about some other girl because I am smart just need help like evryone hear
WITH AREA??? HAVE YOU *NEVER* HEARD OF A CALCULATOR????

valkraybff:

yes I have jerkbag

CrazyHotGhost:

@valkraybff wrote:
yes I have jerkbag
jerkbag

CrazyHotGhost:

i sob.

xXDerpyPugXx:

But do you understand how to answer the problem now?

valkraybff:

no I do not but that ok

Extrinix:

I can't explain it any more than that, multiply.

xXDerpyPugXx:

7 is scale factor, you multiply the scale factor by your other measurements, 1.5 ft and 4 ft

valkraybff:

got it thank

CrazyHotGhost:

Can't be dumbed down no more omg-

valkraybff:

what wrong did you never learn manner

sxdsouls:

what

CrazyHotGhost:

Nah I learned manners but one can only handle so much stupidity in this world

valkraybff:

I not stupid just need help

sxdsouls:

stop being rude you dont know what the user could be going through

Extrinix:

Alright y'all, out of the post, this is not needed anymore lol

CrazyHotGhost:

@extrinix wrote:
Alright y'all, out of the post, this is not needed anymore lol
Alright I got u

CrazyHotGhost:

I'm done.